# Inverse problem for semidiscrete stochastic parabolic operators in arbitrary dimensions

### v1

@@ 2. Semidiscrete inverse random source problem
Compared to EstimateCarlemanStochasticControl and other works on parabolic operators, the weight function here has been modified, particularly in its temporal component; however, the structure $s(t)\varphi(x)$ is preserved. Moreover, the functions satisfy $|\theta(t)|\leq C$ and $|\theta_{t}(t)|\leq C$ for some constant $C>0$ and for all $t\in[0,T]$. As a result, the estimates in Appendix A are obtained, and most of the cross-product terms can be treated in a manner analogous to EstimateCarlemanStochasticControl , slightly simplifying the computations for the new Carleman estimate in this article.
@@
$\begin{cases}d\tilde{w}-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}\tilde{w})=\left(\sum_{i\in\llbracket 1,n\rrbracket}a_{1i}A_{i}D_{i}(\tilde{w})+a_{2}\tilde{w}\right)\,dt+\tilde{g}\,dB(t),&\text{in}\,\mathcal{M}\times(0,T),\\ \tilde{w}=0,&\text{on}\,\partial\mathcal{M}\times(0,T),\\ \tilde{w}(0)=0,\quad&\text{in}\,\mathcal{M}.\end{cases}$
Applying Theorem 1.6 to $\tilde{w}$, we derive the following estimate.
@@
$\begin{split}&J(\tilde{w})+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|\tilde{w}|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|\tilde{g}|^{2}dt\\ &\leq\,C_{s_{0},\varepsilon}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}\left|\sum_{i=1}^{n}a_{1i}A_{i}D_{i}\tilde{w}+a_{2}\tilde{w}\right|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}\tilde{g}|^{2}dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}\tilde{w}|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|\tilde{w}|^{2}\right|_{t=T}\right).\end{split}$
@@
where $J(w)$ is defined as $\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}s\varphi e^{2s\theta\varphi}\,|D_{i}w|^{2}\,dt+\mathbb{E}\int_{Q}s\varphi e^{2s\theta\varphi}|A_{i}D_{i}w|^{2}\,dt$. By applying Young’s inequality for the first integral and using (1.6) for the second term on the right-hand side of the previous inequality, we can rewrite the inequality above as
@@
(2.4) $\begin{split}&J(\tilde{w})+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|\tilde{w}|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|\tilde{g}|^{2}\,dt\\ &\leq C_{s_{0},\varepsilon}\left(\sum_{i=1}^{n}\mathbb{E}\int_{Q}\,e^{2s\varphi}|a_{1i}A_{i}D_{i}(\tilde{w})|^{2}\,dt+\mathbb{E}\int_{Q}\,e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|A_{i}\tilde{g}|^{2}\,dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}\tilde{w}|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|\tilde{w}|^{2}\right|_{t=T}\right).\end{split}$
@@
Observe that, by applying Hölder’s inequality with $\frac{1}{r}+\frac{1}{q}=1$ to the second term on the right-hand side, we obtain
$\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt\leq\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}|a_{2}|^{2r}\,\right)^{1/r}\left(\int_{\mathcal{M}}|e^{s\varphi}\tilde{w}|^{2q}\right)^{1/q}dt.$
Applying the Sobolev-type inequality stated in A.6 with $p=2$ and $p^{\ast}=2q$, we obtain the following.
$\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt\leq\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}|a_{2}|^{2r}\,\right)^{2/r}\left(\int_{\mathcal{M}}|e^{s\varphi}\tilde{w}|^{2}+\sum_{i=1}^{n}\int_{\mathcal{M}}|D_{i}(e^{s\varphi}\tilde{w})|^{2}\right)dt,$
@@
where we can see that choosing $n^{\ast}=2r$ satisfies condition (2.6). Taking into account (1.13) and the triangle inequality, from the above inequality, we deduce
$\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt\leq C\|a_{2}\|_{L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))}$ $\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}|e^{s\varphi}\tilde{w}|^{2}\right$.
$\left.+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}e^{s\varphi}|^{2}|A_{i}\tilde{w}|^{2}+|A_{i}e^{s\varphi}|^{2}|D_{i}\tilde{w}|^{2}\right)dt.$
@@
Notice that thanks to (1.14), for $i=1,..,n$ and $u\in C(\mathcal{M})$ we can obtain the following inequality.
@@
$\int_{\mathcal{M}_{i}^{\ast}}|D_{i}e^{s\varphi}|^{2}|A_{i}\tilde{w}|^{2}\leq\int_{\mathcal{M}_{i}^{\ast}}|D_{i}e^{s\varphi}|^{2}A_{i}(|\tilde{w}|^{2})\leq\int_{\mathcal{M}}A_{i}(|D_{i}e^{s\varphi}|^{2})|\tilde{w}|^{2}$
Therefore,
$\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt\leq C\|a_{2}\|_{L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))}$ $\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}\left(e^{2s\varphi}+\sum_{i=1}^{n}A_{i}(|D_{i}e^{s\varphi}|^{2})\right)|\tilde{w}|^{2}\right$.
$\left.+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}|A_{i}e^{s\varphi}|^{2}|D_{i}\tilde{w}|^{2}\right)dt.$
@@
Finally, using the bounds $A_{i}(|D_{i}(e^{s\varphi})|^{2})\leq\lambda^{2}s^{2}\varphi^{2}(\partial_{i}\psi)^{2}e^{2s\varphi}+e^{2s\varphi}\mathcal{O}_{\lambda}((sh)^{2})$ and $|A_{i}(e^{s\varphi})|^{2}\leq e^{2s\varphi}+e^{2s\varphi}\mathcal{O}_{\lambda}((sh)^{2})$, we conclude that
@@
(2.6) $\begin{split}&\mathbb{E}\int_{Q}e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt\leq C\|a_{2}\|_{L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))}\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}\left(1+\lambda^{2}s^{2}\varphi^{2}|\nabla\psi|^{2}\right)e^{2s\varphi}|\tilde{w}|^{2}\right.\\ &\left.+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}e^{2s\varphi}|D_{i}\tilde{w}|^{2}\right)dt+\mathcal{O}_{\lambda}((sh)^{2})\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}e^{2s\varphi}|\tilde{w}|^{2}+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}e^{2s\varphi}|D_{i}\tilde{w}|^{2}\right)\,dt.\end{split}$
@@
On the other hand, applying (2.5) and performing integration by parts with respect to the averaging operator (1.12), we obtain the following.
$\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|A_{i}\tilde{g}|^{2}\,dt\leq\mathbb{E}\int_{Q_{i}^{\ast}}e^{2s\varphi}A_{i}(|g|^{2})\,dt=\mathbb{E}\int_{Q}A_{i}(e^{2s\varphi})|\tilde{g}|^{2}\,dt+\frac{h}{2}\mathbb{E}\int_{\partial_{i}Q}t_{r}^{i}(e^{2s\varphi})|\tilde{g}|^{2}\,dt.$
Now, considering $\mathbb{E}\int_{\partial_{i}Q}|\tilde{g}|^{2}\,dt=\mathbb{E}\int_{\partial_{i}Q}|d\tilde{w}|^{2}=0$ and $e^{-2s\varphi}A_{i}(e^{2s\varphi})=1+\mathcal{O}((sh)^{2})$, where we used $\tilde{w}=0$ in $\partial_{i}Q$ and Proposition 2.4 in AA:perez:2024 , respectively, we obtain the following.
$\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|A_{i}\tilde{g}|^{2}\,dt\leq\mathbb{E}\int_{Q}\left(e^{2s\varphi}+\mathcal{O}_{\lambda}((sh)^{2})e^{2s\varphi}\right)|\tilde{g}|^{2}\,dt.$
@@
Finally, recording that $s(t)\triangleq\tau\theta(t)$, using the above inequality in the third term on the right-hand side of (2.4) and letting $\tau\geq C$, we can absorb the initial three terms on the right-hand side of the same inequality (2.4). Therefore, we can achieve the following inequality:
@@
$\begin{split}&J(\tilde{w})+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|\tilde{w}|^{2}\,dt+\sum_{i,j\in\llbracket 1,n\rrbracket}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}w|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|\tilde{g}|^{2}\,dt\\ &\leq C\left(\sum_{i\in\llbracket 1,n\rrbracket}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}st_{r}^{i}(e^{2s\varphi}|D_{i}\tilde{w}|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|\tilde{w}|^{2}\right|_{t=T}\right).\end{split}$
From the above, we deduce (1.7) and the proof of the theorem is completed.
3. Semidiscrete Cauchy problem.

### v2

@@ 2. Semi-discrete inverse random source problem
Compared to Lecaros et al. (2026) and other works on parabolic operators, the weight function here has been modified, particularly in its temporal component; however, the structure $s(t)\varphi(x)$ is preserved. Moreover, the functions satisfy $|\theta(t)|\leq C$ and $|\theta_{t}(t)|\leq C$ for some constant $C>0$ and for all $t\in[0,T]$. As a result, the estimates in Boyer et al. (2010b); Pérez (2024) holds, and most of the cross-product terms can be treated in a manner analogous to Lecaros et al. (2026), slightly simplifying the computations for the new Carleman estimate in this article.
@@
$\begin{cases}d\tilde{w}-\sum_{i=1}^{n}D_{i}(\gamma_{i}D_{i}\tilde{w})=\left({\sum_{i=1}^{n}}a_{1i}A_{i}D_{i}(\tilde{w})+a_{2}\tilde{w}\right)\,dt+\tilde{g}\,dB(t),&\text{in}\,Q,\\ \tilde{w}=0,\quad\text{on}\,\partial Q,\quad\quad\quad\left.\tilde{w}\right|_{t=0}=0,\quad\text{in}\,\mathcal{M}.\end{cases}$
Applying Theorem 1.5 to $\tilde{w}$, we derive the following estimate
@@
$\begin{split}&J(\tilde{w})+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|\tilde{w}|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}\tilde{w}|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|\tilde{g}|^{2}dt\\ &\leq\,\mathcal{C}_{\tau_{0},\varepsilon_{0}}\left(\mathbb{E}\int_{Q}\,e^{2s\varphi}\left|\sum_{i=1}^{n}a_{1i}A_{i}D_{i}\tilde{w}+a_{2}\tilde{w}\right|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|D_{i}\tilde{g}|^{2}dt\right.\\ &\left.+\int_{0}^{T}\int_{G_{0}\cap\mathcal{M}}s^{3}\varphi^{3}e^{2s\varphi}|\tilde{w}|^{2}\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|\tilde{w}|^{2}\right|_{t=T}\right).\end{split}$
@@
By applying Young’s inequality for the first integral and using (1.6) for the second term on the right-hand side of the previous inequality, we can rewrite the inequality above as
@@
(2.2) $\begin{split}&J(\tilde{w})+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|\tilde{w}|^{2}\,dt+\sum_{i,j=1}^{n}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}\tilde{w}|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|\tilde{g}|^{2}\,dt\\ &\leq\mathcal{C}_{\tau_{0},\varepsilon_{0}}\left(\sum_{i=1}^{n}\mathbb{E}\int_{Q}\,e^{2s\varphi}|a_{1i}A_{i}D_{i}(\tilde{w})|^{2}\,dt+\mathbb{E}\int_{Q}\,e^{2s\varphi}|a_{2}\tilde{w}|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|A_{i}\tilde{g}|^{2}\,dt\right.\\ &\left.+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}st_{r}^{i}(e^{2s\varphi}|D_{i}\tilde{w}|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|\tilde{w}|^{2}\right|_{t=T}\right).\end{split}$
Our next objective is to estimate the second term on the right-hand side of the above equation, which we consider as $I$.
@@
By applying Hölder’s inequality with $\frac{1}{r}+\frac{1}{q}=1$ to the second term on the right-hand side, we obtain
$I\leq\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}|a_{2}|^{2r}\,\right)^{1/r}\left(\int_{\mathcal{M}}|e^{s\varphi}\tilde{w}|^{2q}\right)^{1/q}dt.$
Applying (Lecaros et al., 2026, Theorem 1.5) to the second term from the right-hand side above with $p=2$ and $p^{\ast}=2q$, we obtain the following.
$I\leq\mathcal{C}\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}|a_{2}|^{2r}\,\right)^{2/2r}\left(\int_{\mathcal{M}}|e^{s\varphi}\tilde{w}|^{2}+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}(e^{s\varphi}\tilde{w})|^{2}\right)dt,$
@@
where $\mathcal{C}$ is a constant independent of $h$ and if $n>2$ thus $2r=n$ or if $n=2$ thus $2r\in(2,+\infty)$. Taking into account (1.14) and the triangle inequality, from the above inequality, we deduce
$I\leq\mathcal{C}\|a_{2}\|^{2}_{L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{2r}(\mathcal{M}))}\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}|e^{s\varphi}\tilde{w}|^{2}+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}|D_{i}e^{s\varphi}|^{2}|A_{i}\tilde{w}|^{2}+|A_{i}e^{s\varphi}|^{2}|D_{i}\tilde{w}|^{2}\right)dt.$
@@
Notice that thanks to (1.15), for $i=1,\ldots,n$ and $u\in C(\mathcal{M})$ we obtain the following inequality.
@@
$\int_{\mathcal{M}_{i}^{\ast}}|D_{i}e^{s\varphi}|^{2}|A_{i}\tilde{w}|^{2}\leq\int_{\mathcal{M}_{i}^{\ast}}|D_{i}e^{s\varphi}|^{2}A_{i}(|\tilde{w}|^{2})=\int_{\mathcal{M}}A_{i}(|D_{i}e^{s\varphi}|^{2})|\tilde{w}|^{2}$
Therefore, we can assert that
$I\leq\mathcal{C}\|a_{2}\|^{2}_{L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{2r}(\mathcal{M}))}\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}\left(e^{2s\varphi}+\sum_{i=1}^{n}A_{i}(|D_{i}e^{s\varphi}|^{2})\right)\right$. $|\tilde{w}|^{2}$
$+\sum_{i=1}^{n}$ $\left.\int_{\mathcal{M}_{i}^{\ast}}|A_{i}e^{s\varphi}|^{2}|D_{i}\tilde{w}|^{2}\right)dt.$
@@
Finally, taking $n^{*}\geq 2r$ and using the bounds $A_{i}(|D_{i}(e^{s\varphi})|^{2})\leq\lambda^{2}s^{2}\varphi^{2}(\partial_{i}\psi)^{2}e^{2s\varphi}+e^{2s\varphi}\mathcal{O}_{\lambda}((sh)^{2})$ and $|A_{i}(e^{s\varphi})|^{2}\leq e^{2s\varphi}+e^{2s\varphi}\mathcal{O}_{\lambda}((sh)^{2})$, we conclude that
@@
(2.4) $\begin{split}I\leq\mathcal{C}\|a_{2}\|^{2}_{L^{\infty}_{\mathcal{F}}(0,T;L_{h}^{n^{\ast}}(\mathcal{M}))}\mathbb{E}\int_{0}^{T}&\left(\int_{\mathcal{M}}\left(1+\lambda^{2}s^{2}\varphi^{2}|\nabla\psi|^{2}\right)e^{2s\varphi}|\tilde{w}|^{2}+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}e^{2s\varphi}|D_{i}\tilde{w}|^{2}\right)\,dt\\ &+\mathcal{O}_{\lambda}((sh)^{2})\mathbb{E}\int_{0}^{T}\left(\int_{\mathcal{M}}e^{2s\varphi}|\tilde{w}|^{2}+\sum_{i=1}^{n}\int_{\mathcal{M}_{i}^{\ast}}e^{2s\varphi}|D_{i}\tilde{w}|^{2}\right)\,dt,\end{split}$
where we check at once that $n^{\ast}$ satisfies condition (2.4).
@@
On the other hand, let us focus on the term $\mathbb{E}\int_{Q_{i}^{\ast}}|A_{i}\tilde{g}|^{2}\,dt$ in (2.2). Applying (2.3) and performing integration by parts with respect to the averaging operator (1.13), we obtain the following.
(2.5) $\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|A_{i}\tilde{g}|^{2}\,dt\leq\mathbb{E}\int_{Q_{i}^{\ast}}e^{2s\varphi}A_{i}(|\tilde{g}|^{2})\,dt=\mathbb{E}\int_{Q}A_{i}(e^{2s\varphi})|\tilde{g}|^{2}\,dt+\frac{h}{2}\mathbb{E}\int_{\partial_{i}Q}t_{r}^{i}(e^{2s\varphi})|\tilde{g}|^{2}\,dt.$
To deal with the last term in the above inequality, consider $x_{i}=e_{i}\cdot x$, we can assert that
$\int_{\mathcal{M}}(2x_{i}-1)D_{i}(e^{2s\varphi})|\tilde{g}|^{2}=-\int_{\mathcal{M}_{i}^{*}}2e^{2s\varphi}A_{i}(|\tilde{g}|^{2})-\int_{\mathcal{M}_{i}^{*}}A_{i}(2x_{i}-1)e^{2s\varphi}D_{i}(|\tilde{g}|^{2})$
$+\int_{\partial_{i}\mathcal{M}}(2x_{i}-1)\nu_{i}t_{r}^{i}(e^{2s\varphi})|\tilde{g}|^{2}.$
[…]
$\int_{\partial_{i}\mathcal{M}}(2x_{i}-1)\nu_{i}$ $t_{r}^{i}(e^{2s\varphi})|\tilde{g}|^{2}\leq\int_{\mathcal{M}}(2x_{i}-1)D_{i}(e^{2s\varphi})|\tilde{g}|^{2}+\int_{\mathcal{M}_{i}^{\ast}}2e^{2s\varphi}A_{i}(|\tilde{g}|^{2})$
$+\int_{\mathcal{M}_{i}^{*}}A_{i}(2x_{i}-1)e^{2s\varphi}|A_{i}\tilde{g}|^{2}+\int_{\mathcal{M}_{i}}A_{i}(2x_{i}-1)e^{2s\varphi}|D_{i}\tilde{g}|^{2}.$
Applying again the condition (1.6) and the fact $|A_{i}\tilde{g}|^{2}\leq A_{i}(|\tilde{g}|^{2})$, we conclude that
$\int_{\partial_{i}\mathcal{M}}(2x_{i}-1)\nu_{i}t_{r}^{i}(e^{2s\varphi})|\tilde{g}|^{2}\leq$ $\int_{\mathcal{M}}(2x_{i}-1)D_{i}(e^{2s\varphi})|\tilde{g}|^{2}$
$+2\int_{\mathcal{M}_{i}^{\ast}}(1+A_{i}(2x_{i}-1))e^{2s\varphi}A_{i}(|\tilde{g}|^{2}).$
Moreover, using that $|1+A_{i}(2x_{i}-1)|\leq 2$ for all $x_{i}$ and integration by parts with respect to average operator, it follows that
$\int_{\partial_{i}\mathcal{M}}(2x_{i}-1)\nu_{i}t_{r}^{i}(e^{2s\varphi})|\tilde{g}|^{2}\leq\int_{\mathcal{M}}(2x_{i}-1)D_{i}(e^{2s\varphi})|\tilde{g}|^{2}+4\int_{\mathcal{M}}A_{i}(e^{2s\varphi})|\tilde{g}|^{2}+2h\int_{\partial_{i}\mathcal{M}}t_{r}^{i}(e^{2s\varphi})|\tilde{g}|^{2}$
Taking $h_{1}\leq\frac{1}{4}$ and noting that $(2x_{i}-1)\nu_{i}=1$ for $x_{i}\in\partial_{i}\mathcal{M}$, we have
$\frac{1}{2}\int_{\partial_{i}\mathcal{M}}t_{r}^{i}(e^{2s\varphi})|\tilde{g}|^{2}\leq\int_{\mathcal{M}}(2x_{i}-1)D_{i}(e^{2s\varphi})|\tilde{g}|^{2}+4\int_{\mathcal{M}}A_{i}(e^{2s\varphi})|\tilde{g}|^{2}$
for all $0<h<h_{1}$.
Combining the above equation and (2.5) with $e^{-2s\varphi}D_{i}(e^{2s\varphi})=s\lambda\varphi\partial_{i}\psi+\mathcal{O}_{\lambda}((sh)^{2})$ and $e^{-2s\varphi}A_{i}(e^{2s\varphi})=1+\mathcal{O}_{\lambda}((sh)^{2})$, due to Proposition 2.4 in Pérez (2024), respectively, we obtain
$\mathbb{E}\int_{Q_{i}^{\ast}}\,e^{2s\varphi}|A_{i}\tilde{g}|^{2}\,dt\leq\mathbb{E}\int_{Q}\left(1+\mathcal{O}_{\lambda}(sh)+\mathcal{O}_{\lambda}(h)+\mathcal{O}_{\lambda}((sh)^{2})\right)e^{2s\varphi}|\tilde{g}|^{2}\,dt.$
for all $0<h<\min\{h_{0},h_{1}\}$.
@@
Finally, recalling that $s(t):=\tau\theta(t)$, using the above inequality in the third term on the right-hand side of (2.2), the estimate of $I$ and letting $\tau\geq C$, we can absorb the initial three terms on the right-hand side of the same inequality (2.2). Therefore, we can achieve the following inequality,
@@
$\begin{split}&J(\tilde{w})+\mathbb{E}\int_{Q}s^{3}\lambda^{4}_{1}\varphi^{3}\,e^{2s\varphi}|\tilde{w}|^{2}\,dt+{\sum_{i,j=1}^{n}}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1}e^{2s\varphi}|D_{ij}^{2}\tilde{w}|^{2}\,dt+\mathbb{E}\int_{Q}s\lambda_{1}^{2}\varphi\,e^{2s\varphi}|\tilde{g}|^{2}\,dt\\ &\leq\mathcal{C}\left(\int_{0}^{T}\int_{G_{0}\cap\mathcal{M}}s^{3}\varphi e^{2s\varphi}|\tilde{w}|\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}\,e^{2s\varphi}|\tilde{w}|^{2}\right|_{t=T}\right).\end{split}$
After dropping the positive terms, from the left-hand side of above inequality and recalling the definition of the operator $\Lambda_{1}$, we deduce (1.7) and the proof of the theorem is completed.
3. Semi-discrete Cauchy problem.